Please help me find the fifth term of (x+y)^12 and the last term of (2x-y)^11 could you help me understand this too? I dont even get the question...let alone how to find the answer. Thanks!
Answer by stanbon(75887)

Please help me find the fifth term of (x+y)^12 and the last term of (2x-y)^11 could you help me understand this too? I dont even get the question...let alone how to find the answer. Thanks!
The 5th terms is 12C4x^4(-y)^8= 495x^4y^8
The last term is the 12th terms = 11C0x^0(-y)^11=-y^11
Cheers,
Stan H.